Circular No. 7774

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 X1 (LINEAR) R. Huber, Lincoln Laboratory, Massachusetts Institute of Technology,

reports the discovery by LINEAR of a comet on images taken on Dec. 13 (discovery observation given below) that show a bright core surrounded by a diffuse coma (comet’s magnitude given as 16.5–17.0) and exhibiting a tail at least 95′′ long in p.a. 295^{o}. CCD observations on Dec. 14.1 UT by R. Stoss and P. Geffert (Starkenburg Sternwarte, 0.45-m f/4.4 reflector) reveal a well-condensed coma with a thin tail 5′ long in p.a. 300^{o} (m_{2} = 16.5). CCD images taken in twilight and poor seeing by A. C. Gilmore (Mount John, 1.0-m f/7.7 reflector) on Dec. 14.6 show the object as diffuse (m_{1} = 14.6–15.1) with a broad, faint fan tail 1′ long in p.a. ∼ 315^{o}.

2001 UT  α_{2000}  δ_{2000}
Dec. 13.44178  12^{h}15^{m}50.^{s}63  −7^{o}57′00′′.9

Further astrometric observations and the following preliminary parabolic orbital elements appear on MPEC 2001-X55:

T = 2001 Dec. 30.277 TT  ω = 196.^{o}345
Ω = 336.252 }2000.0
q = 1.73851 AU  i = 113.293
2001TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 7 12^{h}21.^{m}06 − 4^{o}28.^{′}2 1.873 1.765 68.^{o}3 31.^{o}3 14.8
12 12 17.19 − 7 08.0 1.763 1.755 73.3 32.5 14.7
17 12 11.97 −10 03.9 1.654 1.747 78.4 33.5 14.5
22 12 05.04 −13 18.8 1.548 1.742 83.8 34.2 14.4
27 11 55.92 −16 55.0 1.446 1.739 89.3 34.4 14.2

SUPERNOVA 2001ie IN UGC 5542 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370-750 nm) of SN 2001ie (cf. IAUC 7771), obtained by W. Brown on Dec. 12.48 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum light. Adopting the NASA/IPAC Extragalactic Database recession velocity of 9215 km/s for the host galaxy, the supernova expansion velocity is ∼ 11 500 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) was then 3 ± 2 days after maximum light.

2001 December 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green